U=x^2y^2+xy Derive the Marshallian demands and indirect utility function

U=x^2y^2+xy

Derive the Marshallian demands and indirect utility function
Answer #1

Marshallian Demand

AnS U= x²y2 tay Marshallian Demands Budget constraint = M - 3 P = Price of x Pix + P2 g = Pax + P2 get constraint e = price o= 24g + g - 0 2x²y + x-1 P = 0 2x²y tx = dpz – 2x2y + x d = 2 Equeeting value of d- y = x Putting value of y in egr & м= f, яM= P, R tk (2x) M = Pint P a M = 2 pix ② Putting value of x* in ear y = x Marshallian Demand **

Indirect Utility Function V(P, M)

For Indirect utility substituting value of xt and yt in U U = x²y2 + xy - IUF 12 M2 UUUUUUUUUUUU866 Indirect utility function

​​​​​